# Decentralized and Uncoordinated Learning of Stable Matchings: A Game-Theoretic Approach

### v1

@@ II A Game-Theoretic Formulation and Preliminary Results
Moreover, given a strategy profile $x=(x_{m},x_{-m})$, we define the payoff of player $m$ by  [^5] Here, $x_{-m}$ refers to the strategy vector of all the players except the $m$th one.
@@
One can interpret $u_{m}(x)$ as the expected utility that player $m$ would receive by successfully getting matched if each man independently proposes to women according to his mixed strategy distribution $x_{m}$. The reason for defining the payoff functions as in (1) is that we want our devised algorithms to be implemented in a fully decentralized and uncoordinated manner among men by relying only on their received feedback (embedded into their payoff functions). While such payoff functions are highly nonlinear, they are essential to eliminate any degrees of coordination among the players. In other words, the cost of devising a fully coordination-free algorithm comes in analyzing more complex payoff functions with higher degrees of nonlinearity. We refer to Section VI for a simpler form of payoff functions with less degree of nonlinearities, which in turn requires stronger information feedback. The following are three properties of the payoff functions (1).
@@
In the remainder of this paper, we will refer to the above noncooperative game as the stable matching game and simply denote it by $\mathcal{G}=(M,\{u_{m}\}_{m\in M},\{\mathcal{X}_{m}\}_{m\in M})$. It is worth noting that although we are interested in obtaining a NE of the stable matching game, which is a complete information one-shot game, our goal is to learn one such NE by repeatedly playing the incomplete information game, in which the true preferences $\mu_{mw}$ (and hence the payoff functions) are not fully known. In fact, the set of pure/mixed NE points of the stable matching game has some interesting connections with the set of stable matchings with known preferences. The following theorem establishes one such result.
@@
As Example 1 suggests, mixed NE points may represent unpredictable bipartite graphs (Remark 1), and it is difficult to make general statements about their structure. However, the following theorem establishes an interesting connection between mixed and pure NE points, implying that obtaining a mixed NE with full support on the women’s side is as satisfactory as obtaining a pure NE.
@@
Let $x$ be any mixed NE in which each woman receives at least one (fractional) proposal. [^6] If this condition does not hold, the theorem statement continues to hold for the induced submarket where each woman has received at least one fractional proposal. For each man $m$, let $w_{m}$ be the least preferred woman among those that he fractionally proposes to them, i.e., $w_{m}=\mathop{\rm argmin}_{w\in W}\{\mu_{mw}:x_{mw}>0\}$. Then $\{(m,w_{m}),m\in M\}$ forms a stable matching (and hence a pure NE) for the stable matching game.
@@
Please see Appendix A-B for the proof.  Q.E.D.
@@
III Logarithmic Regret for Hierarchical Markets
In this section, we consider the stable matching game with unknown preferences and develop a decentralized and uncoordinated algorithm for learning its NE points. In particular, we focus on matching markets with specific structures because without imposing any assumption on the matching market, there are exponential lower bounds for the number of iterations to find a stable matching through better response dynamics [11]. One example of such a structured matching market is the hierarchical matching market, where it was shown in [2] that if men follow a certain decentralized (but coordinated) algorithm, the men’s expected regret (see Definition 3) is at most logarithmic in time but exponential in terms of other parameters such as the number of men $n$. In particular, the authors in [2] raised the question of how to remove the exponential dependence of the regret on $n$. In this section, we show that Algorithm 1 achieves this goal whose regret is logarithmic in time with a polynomial factor in terms of other parameters. Moreover, Algorithm 1 is uncoordinated and, hence, removes the phase-dependent coordination required by the algorithm in [2]. Throughout this section, we impose the following hierarchical assumption on the matching market.
@@ III Logarithmic Regret for Hierarchical Markets
The condition imposed in Assumption 1 is also known as the Sequential Preference Condition (SPC) [14] and is weaker than the $\alpha$-reducible condition considered in [2], which assumes that each submarket has a fixed-pair: A pair $(m,w)$ is called a fixed pair if both $m$ and $w$ prefer each other the most among anyone else in that submarket. [^7] In fact, one can show that a matching market has a unique stable matching if and only if it satisfies $\alpha$-reducible condition [2]. Since Assumption 1 is weaker than the $\alpha$-reducible condition, all our results immediately applies to $\alpha$-reducible markets.
@@
Let us assume that players follow a decentralized and uncoordinated algorithm $\mathcal{A}$ that results in a sequence of proposals $\alpha^{t}=(\alpha^{t}_{m},m\in M),t=1,\ldots,T$. The expected regret of $\mathcal{A}$ equals
@@
which is the expected number of times that the men’s proposals do not form a stable matching.
III-A Algorithm Design and Preliminaries
The algorithm that we propose is an adaptation of the online dual mirror descent for adversarial multi-arm bandit problem (see, e.g., EXP3 Algorithm in [18, Chapter 11]) to the stable matching game, where for simplicity we refer to it as the EXP algorithm. In particular, it uses an entropy regularizer and adaptive stepsize for updating the players’ mixed strategy distributions over time.
@@ III-A Algorithm Design and Preliminaries
Input: A decreasing stepsize sequence $\{\eta^{t}\}$, a mixing sequence $\{\gamma^{t}\}$, and an initial vector $\hat{L}^{0}_{m}=\boldsymbol{0}$.
For $t=1,2,\ldots$, player $m$ independently performs the following steps:
@@
We begin with the following lemma, which shows that the feedback $\hat{v}^{t}_{m}$ received by player $m$ at time $t$ is conditionally an unbiased and bounded estimator of the actual gradient vector $v_{m}^{t}$.
@@
The received feedback is a conditionally unbiassed and bounded estimate of the gradients
@@
where the second equality holds because conditioned on $\mathcal{F}^{t-1}$, the random variables $\{\alpha^{t}_{k},\forall k\}$ are mutually independent as they are drawn independently from the given distributions $\{\hat{X}^{t}_{k},\forall k\}$. Similarly,
@@
III-B Performance of Algorithm 1 for Hierarchical Markets
Here, we analyze the performance of Algorithm 1 for hierarchical matching markets. In the following analysis, we assume that a positive lower bound for $c:=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$ is known to all the players so that they can choose their mixing parameters in Algorithm 1. However, as we discuss in Appendix A-G, this condition can be relaxed up to some extent by using an uncoordinated and decentralized UCB gap estimation oracle in which the players use their estimated gaps to tune their mixing parameters. The main result of this section is presented in Theorem 3, which essentially shows that for hierarchical matching markets, Algorithm 1 achieves an expected regret that is logarithmic in $T$ and polynomial in terms of other parameters. However, before delving into the analysis, we first provide an overview of the main steps in the proof.
@@ III-B Performance of Algorithm 1 for Hierarchical Markets
  • Conditioned on the event $\Omega$, Lemma 3 shows that, due to the market structure and payoff functions, as time progresses, more pairs in the market are matched according to the underlying hierarchical order with very high probability. Additionally, the size of such hierarchical nested matchings grows with very high probability after a polynomial number of steps.
@@
Let us consider the stochastic dynamics of Algorithm 1 that are given by
@@
$\hat{X}^{t}_{m}=(1-\gamma^{t})X^{t}_{m}+\frac{\gamma^{t}}{n}\boldsymbol{1},$ (13)
@@
where $c=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$ is a constant parameter. The following lemma provides a high probability bound for the event $\Omega$ to occur under appropriate choice of stepsize and mixing parameter.
@@
where $S^{t-1}_{mw}:=\sum_{\tau=1}^{t-1}\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}\ \forall m,w$. According to Lemma 1, $\{\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\}$ is a martingale difference sequence that satisfies
$|\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})|\leq\frac{n}{\gamma^{\tau}},\ \ \ \ \mathbb{E}[\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}^{2}|\mathcal{F}^{\tau-1}]\leq\frac{n}{\gamma^{\tau}}.$
By taking $u=\frac{n}{\gamma^{t-1}}$, $v^{2}=\sum_{\tau=1}^{t-1}\frac{n}{\gamma^{\tau}}$, and $x=ct$ in Lemma 7 (see Appendix A-A) and noting that the event $\{V^{t}\leq v^{2}\}$ holds with probability 1, we obtain
@@
Assume that each man follows Algorithm 1 with stepsize $\eta^{t}=\frac{1}{\sqrt{t}}$ and mixing parameter $\gamma^{t}=M\frac{\log t}{t}$, where $M=\frac{4n}{c}\log\frac{1}{\delta}$. Let $a_{1}\geq\ldots\geq a_{n}$ be a sequence defined by $a_{1}=\frac{\Delta}{2}$ and $a_{k}=\frac{1}{4}\min_{i\in[k]}\big{\{}\Delta,\mu_{m_{i}w_{i}}\big{\}},\forall k=2,\ldots,n$. Conditioned on the event $\Omega$, there exist $t_{1}\leq\ldots\leq t_{n}=\tilde{O}\big{(}\frac{n^{2}}{c^{2}}\log\frac{1}{\delta}\big{)}$, such that for any $t\geq t_{k}$ and any $w\neq w_{k}$, we have
$\hat{X}^{t}_{m_{k}w_{k}}\geq\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}+\frac{\gamma^{t}}{n},\ \ \ \hat{X}^{t}_{m_{k}w}\leq\frac{(1-\gamma^{t})e^{-\eta^{t}a_{k}t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}+\frac{\gamma^{t}}{n}.$ (21)
@@
where the third inequality holds because $c\leq\frac{a_{1}}{4}$, and the last inequality holds because $t\geq t_{1}$. Moreover,
@@
Combining relations (24) and (27) and using the fact that $\hat{X}^{t}_{mw}=(1-\gamma^{t})X^{t}_{mw}+\frac{\gamma^{t}}{n}$, for any $t\geq t_{1}$ and $w\neq w_{1}$, we have
$\hat{X}^{t}_{m_{1}w_{1}}\geq\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{1}t}}+\frac{\gamma^{t}}{n},\ \ \ \ \ \ \ \ \hat{X}^{t}_{m_{1}w}\leq\frac{(1-\gamma^{t})e^{-\eta^{t}a_{1}t}}{1+(n-1)e^{-\eta^{t}a_{1}t}}+\frac{\gamma^{t}}{n}.$ (28)
@@
$Y^{t-1}_{m_{k}w_{k}}$ $-Y^{t-1}_{m_{k}w}=\sum_{\tau=1}^{t-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w_{i}}(\hat{X}^{\tau})\big{)}$
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\prod_{\ell=1}^{k-1}(1-\hat{X}^{\tau}_{m_{\ell}w_{k}})-\mu_{m_{k}w_{i}}(1-\hat{X}^{\tau}_{m_{i}w_{i}})\Big{)}$ (31)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\big{(}1-\sum_{\ell=1}^{k-1}(\frac{\gamma^{t}}{n}+\frac{(1-\gamma^{t})e^{-\eta^{t}a_{\ell}t}}{1+(n-1)e^{-\eta^{t}a_{\ell}t}})\big{)}-\mu_{m_{k}w_{i}}\big{(}1-\frac{\gamma^{t}}{n}-\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{i}t}}\big{)}\Big{)}$ (32)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}-\gamma^{t}-\sum_{\ell=1}^{k-1}\frac{(1-\gamma^{t})e^{-\eta^{t}a_{\ell}t}}{1+(n-1)e^{-\eta^{t}a_{\ell}t}}-\mu_{m_{k}w_{i}}\frac{(n-1)e^{-\eta^{t}a_{i}t}+\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{i}t}}\Big{)}$ (33)
$\geq\sum_{\tau=1}^{t-1}\Big{(}\mu_{m_{k}w_{k}}-\gamma^{t}-\sum_{\ell=1}^{k-1}e^{-\eta^{t}a_{\ell}t}-\mu_{m_{k}w_{i}}(n-1)e^{-\eta^{t}a_{i}t}-\mu_{m_{k}w_{i}}\gamma^{t}\Big{)}$ (34)
$\geq\sum_{\tau=1}^{t-1}\big{(}\mu_{m_{k}w_{k}}-2\gamma^{t}-2(n-1)e^{-\eta^{t}a_{k-1}t}\big{)},$ (35)
where the second inequality holds by the induction hypothesis for $t\geq t_{k-1}$, and the last inequality holds because $a_{1}\geq\ldots\geq a_{k-1}$. Thus, for any time step $t\geq t_{k-1}$ that satisfies $e^{-\eta^{t}a_{k-1}t}\leq\gamma^{t}$ and $\frac{t-2}{\sum_{\tau=1}^{t-1}\gamma^{\tau}}\geq\frac{4n}{\mu_{m_{k}w_{k}}}$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}$ $\geq\sum_{\tau=1}^{t-1}\big{(}\mu_{m_{k}w_{k}}-2n\gamma^{t}\big{)}=\mu_{m_{k}w_{k}}(t-1)-2n\sum_{\tau=1}^{t-1}\gamma^{t}\geq\frac{\mu_{m_{k}w_{k}}}{2}t.$ (36)

### v2

@@ III-A Stable Matching Game
Let $x_{-m}$ denote the strategy vector of all the players except the $m$th one. Given a strategy profile $x=(x_{m},x_{-m})$, we define the payoff of player $m$ by
@@
One can interpret $u_{m}(x)$ as the expected utility that player $m$ would receive by successfully getting matched if each man independently proposes to women according to his mixed strategy distribution $x_{m}$. The reason for defining the payoff functions as in (1) is that we want our devised algorithms to be implemented in a fully decentralized and uncoordinated manner among men by relying only on their received feedback (embedded into their payoff functions). While such payoff functions are highly nonlinear, they are essential to eliminate any degrees of coordination among the players. In other words, the cost of devising a fully coordination-free algorithm comes in analyzing more complex payoff functions with higher degrees of nonlinearity. [^1] We refer to Appendix B for a simpler form of payoff functions with less degree of nonlinearities, which in turn requires stronger information feedback. The following are three properties of the payoff functions given in (1).
@@
In the remainder of this paper, we will refer to the above noncooperative game as the stable matching game and denote it by $\mathcal{G}=(M,\{u_{m}\}_{m\in M},\{\mathcal{X}_{m}\}_{m\in M})$. It is important to note that although our interest is in obtaining a NE of the stable matching game, which is a complete information one-shot game, our goal is to learn such a NE by repeatedly playing the incomplete information game, where the true preferences $\mu_{mw}$ (and hence the payoff functions) are not fully known.
III-B Nash Equilibrium Characterization
In this part, we show that the set of pure and mixed NE points of the stable matching game has interesting connections with the set of stable matchings when preferences are known. The following theorem establishes one such result.
@@ III-B Nash Equilibrium Characterization
As Example 1 suggests, mixed NE points may correspond to unpredictable bipartite graphs (Remark 1), making it difficult to make general statements about their structure. However, the following theorem establishes an interesting connection between mixed and pure NE points, implying that obtaining a mixed NE with full support on the women’s side is as satisfactory as obtaining a pure NE.
@@
Let $x$ be any mixed NE in which each woman receives at least one (fractional) proposal. For each man $m$, let $w_{m}$ be the least preferred woman among those that he fractionally proposes to them, i.e., $w_{m}=\mathop{\rm argmin}_{w\in W}\{\mu_{mw}:x_{mw}>0\}$. Then $\{(m,w_{m}),m\in M\}$ forms a stable matching (and hence a pure NE) for the stable matching game.
@@
$\prod_{k>_{w_{1}}m}(1-x_{kw_{1}})=\prod_{k>_{w_{2}}m}(1-x_{kw_{2}})=1,$
which together with the above argument implies that
Remark 2
As can be seen from the proof of Theorem 2, if the condition that each woman receives at least one proposal does not hold, the theorem still applies to the induced submarket in which each woman has received at least one fractional proposal.
@@
IV Logarithmic Regret for Hierarchical Markets
In this section, we consider the stable matching game with unknown preferences and develop a decentralized and uncoordinated algorithm for learning its NE points. In particular, we focus on matching markets with specific structures because without imposing any assumption on the matching market, there are exponential lower bounds for the number of iterations to find a stable matching through better response dynamics [11]. One example of such a structured matching market is the hierarchical matching market, where it was shown in [2] that if men follow a certain decentralized (but coordinated) algorithm, the men’s expected regret (see Definition 3) is at most logarithmic in time but exponential in terms of other parameters such as the number of men $n$. In this section, we show that a much simpler algorithm (Algorithm 1) also achieves logarithmic regret in a fully uncoordinated fashion, hence, removing the phase-dependent coordination required by the algorithm in [2]. In particular, our analysis provides a cleaner characterization of the other constants involved in the regret bound. Throughout this section, we impose the following hierarchical assumption on the matching market.
@@ IV Logarithmic Regret for Hierarchical Markets
The condition imposed in Assumption 1 is also known as the Sequential Preference Condition (SPC) [18] and is weaker than the $\alpha$-reducible condition considered in [2], which assumes that each submarket has a fixed-pair: A pair $(m,w)$ is called a fixed pair if both $m$ and $w$ prefer each other the most among anyone else in that submarket. In fact, one can show that a matching market has a unique stable matching if and only if it satisfies $\alpha$-reducible condition [2]. Since Assumption 1 is weaker than the $\alpha$-reducible condition, all our results immediately applies to $\alpha$-reducible markets.
@@
Let us assume that players follow a decentralized and uncoordinated algorithm $\mathcal{A}$ that results in a sequence of proposals $\alpha^{t}=(\alpha^{t}_{m},m\in M),t=1,\ldots,T$. The expected regret of algorithm $\mathcal{A}$ is defined by
@@
where $\boldsymbol{1}_{\{\cdot\}}$ denotes the indicator function. In other words, $R(T)$ is the expected number of times that the men’s proposals do not form a stable matching.
IV-A Algorithm Design and Preliminaries
The algorithm that we propose is an adaptation of the online dual mirror descent for adversarial multi-arm bandit problem (see, e.g., EXP3 Algorithm in [19, Chapter 11]) to the stable matching game, where for simplicity we refer to it as the EXP algorithm. In particular, it uses an entropy regularizer and adaptive stepsize for updating the players’ mixed strategy distributions over time.
@@ IV-A Algorithm Design and Preliminaries
Input: A decreasing stepsize sequence $\{\eta^{t}\}$, a mixing sequence $\{\gamma^{t}\}$, and an initial vector $\hat{L}^{0}_{m}=\boldsymbol{0}$. For $t=1,2,\ldots$, player $m$ independently performs the following steps:
@@
We begin with the following lemma, which shows that the feedback $\hat{v}^{t}_{m}$ received by player $m$ at time $t$ is conditionally an unbiased and bounded estimator of the actual gradient vector $v_{m}^{t}$. The proof of this lemma follows from standard analysis of the multi-arm bandit setting [17]. However, for the sake of completeness, we provide a short proof here.
@@
The received feedback is a conditionally unbiassed and bounded estimate of the gradients:
@@
where the second equality holds because conditioned on $\mathcal{F}^{t-1}$, the random variables $\{\alpha^{t}_{k},\forall k\}$ are mutually independent as they are drawn independently from the given distributions $\{\hat{X}^{t}_{k},\forall k\}$. Similarly, we can bound the expected second moment as
@@
IV-B Performance of Algorithm 1 for Hierarchical Markets
Here, we analyze the performance of Algorithm 1 for hierarchical matching markets. In the following analysis, we assume that a positive lower bound for $c:=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$ is known to all the players so that they can choose their mixing parameters in Algorithm 1. However, as we discuss in Appendix B-E, this condition can be relaxed up to some extent by using an uncoordinated and decentralized UCB gap estimation oracle in which the players use their estimated gaps to tune their mixing parameters. The main result of this section is presented in Theorem 3, which essentially shows that for hierarchical matching markets, Algorithm 1 achieves an expected regret that is logarithmic in $T$ without any coordination. However, before delving into the analysis, we first provide an overview of the main steps in the proof.
@@ IV-B Performance of Algorithm 1 for Hierarchical Markets
  • Conditioned on the event $\Omega$, Lemma 3 shows that, due to the market structure and payoff functions, as time progresses, more pairs in the market are matched according to the underlying hierarchical order with high probability. Additionally, the size of such nested matchings grows with high probability after a fixed number of steps.
@@
Now we are ready to begin our main analysis. Let us consider the stochastic dynamics of Algorithm 1, which for any pair $(m,w)$ are described by
@@
$\hat{X}^{t}_{mw}=(1-\gamma^{t})X^{t}_{mw}+\frac{\gamma^{t}}{n},$ (13)
@@
where $c=\frac{1}{8}\min_{k\in[n]}\{\Delta,\mu_{m_{k}w_{k}}\}$ is a constant. The following lemma provides a high probability bound for the event $\Omega$ to occur under appropriate choice of stepsize and mixing parameter.
@@
where $S^{t-1}_{mw}:=\sum_{\tau=1}^{t-1}\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}\ \forall m,w$. According to Lemma 1, $\{\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\}$ is a martingale difference sequence that almost surely satisfies
$|\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})|\leq\frac{n}{\gamma^{\tau}},$
$\mathbb{E}[\big{(}\hat{v}^{\tau}_{mw}-v_{mw}(\hat{X}^{\tau})\big{)}^{2}|\mathcal{F}^{\tau-1}]\leq\frac{n}{\gamma^{\tau}}.$ (17)
By taking $u=\frac{n}{\gamma^{t-1}}$, $v^{2}=\sum_{\tau=1}^{t-1}\frac{n}{\gamma^{\tau}}$, and $x=ct$ in Lemma 7 (see Appendix B-A) and noting that the event $\{V^{t}\leq v^{2}\}$ holds with probability 1, we obtain
@@
Assume that each man follows Algorithm 1 with stepsize $\eta^{t}=\frac{1}{\sqrt{t}}$ and mixing parameter $\gamma^{t}=M\frac{\log t}{t}$, where $M=\frac{4n}{c}\log\frac{1}{\delta}$. Let $a_{1}\geq\ldots\geq a_{n}$ be a sequence defined by $a_{1}=\frac{\Delta}{2}$ and $a_{k}=\frac{1}{4}\min_{i\in[k]}\big{\{}\Delta,\mu_{m_{i}w_{i}}\big{\}},\forall k=2,\ldots,n$. Conditioned on the event $\Omega$, there exists a sequence of time instances $t_{1}\leq\ldots\leq t_{n}=O\big{(}\frac{nM}{c^{n+1}}\log^{2}(\frac{M}{c})\big{)}$, such that for any $k\in[n]$, $t\geq t_{k}$, and $w\neq w_{k}$, we have
$\hat{X}^{t}_{m_{k}w_{k}}\geq\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}+\frac{\gamma^{t}}{n},$ (22)
$\hat{X}^{t}_{m_{k}w}\leq\frac{(1-\gamma^{t})e^{-\eta^{t}a_{k}t}}{1+(n-1)e^{-\eta^{t}a_{k}t}}+\frac{\gamma^{t}}{n}.$ (23)
@@
where the third inequality holds because $c\leq\frac{a_{1}}{4}$, and the last inequality holds by $t\geq t_{1}$. Moreover,
@@
Combining (25) and (28) and using $\hat{X}^{t}_{mw}=(1-\gamma^{t})X^{t}_{mw}+\frac{\gamma^{t}}{n}$, for any $t\geq t_{1}$ and $w\neq w_{1}$, we have
$\hat{X}^{t}_{m_{1}w_{1}}\geq\frac{1-\gamma^{t}}{1+(n-1)e^{-\eta^{t}a_{1}t}}+\frac{\gamma^{t}}{n},$ (29)
$\hat{X}^{t}_{m_{1}w}\leq\frac{(1-\gamma^{t})e^{-\eta^{t}a_{1}t}}{1+(n-1)e^{-\eta^{t}a_{1}t}}+\frac{\gamma^{t}}{n}.$ (30)
@@
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}=\sum_{\tau=t_{k-1}}^{t-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w_{i}}(\hat{X}^{\tau})\big{)}+\sum_{\tau=1}^{t_{k-1}-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w_{i}}(\hat{X}^{\tau})\big{)}$
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\prod_{\ell=1}^{k-1}(1-\hat{X}^{\tau}_{m_{\ell}w_{k}})-\mu_{m_{k}w_{i}}(1-\hat{X}^{\tau}_{m_{i}w_{i}})\Big{)}+\sum_{\tau=1}^{t_{k-1}-1}\big{(}v_{m_{k}w_{k}}(\hat{X}^{\tau})-v_{m_{k}w_{i}}(\hat{X}^{\tau})\big{)}$ (33)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}\big{(}1-\sum_{\ell=1}^{k-1}(\frac{\gamma^{\tau}}{n}+\frac{(1-\gamma^{\tau})e^{-\eta^{\tau}a_{\ell}\tau}}{1+(n-1)e^{-\eta^{\tau}a_{\ell}\tau}})\big{)}-\mu_{m_{k}w_{i}}\big{(}1-\frac{\gamma^{\tau}}{n}-\frac{1-\gamma^{\tau}}{1+(n-1)e^{-\eta^{\tau}a_{i}\tau}}\big{)}\Big{)}-t_{k-1}$ (34)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}-\gamma^{\tau}-\sum_{\ell=1}^{k-1}\frac{(1-\gamma^{\tau})e^{-\eta^{\tau}a_{\ell}\tau}}{1+(n-1)e^{-\eta^{\tau}a_{\ell}\tau}}-\frac{(n-1)e^{-\eta^{\tau}a_{i}\tau}+\gamma^{\tau}}{1+(n-1)e^{-\eta^{\tau}a_{i}\tau}}\Big{)}-t_{k-1}$ (35)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\Big{(}\mu_{m_{k}w_{k}}-\gamma^{\tau}-\sum_{\ell=1}^{k-1}e^{-\eta^{\tau}a_{\ell}\tau}-(n-1)e^{-\eta^{\tau}a_{i}\tau}-\gamma^{\tau}\Big{)}-t_{k-1}$ (36)
$\geq\sum_{\tau=t_{k-1}}^{t-1}\big{(}\mu_{m_{k}w_{k}}-2\gamma^{\tau}-2ne^{-\eta^{\tau}a_{k-1}\tau}\big{)}-t_{k-1},$ (37)
where the second inequality holds by the induction hypothesis for $\tau\geq t_{k-1}$, and the last inequality holds because $a_{1}\geq\ldots\geq a_{k-1}$. Using the specific choice of stepsize and mixing parameters $\eta^{t}=\frac{1}{\sqrt{t}}$ and $\gamma^{t}=M\frac{\log t}{t}$, for any $t\geq t_{k-1}$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}$ $\geq\sum_{\tau=t_{k-1}}^{t-1}\big{(}\mu_{m_{k}w_{k}}-2\gamma^{\tau}-2ne^{-\eta^{\tau}a_{k-1}\tau}\big{)}-t_{k-1}$
[…]
$\geq\mu_{m_{k}w_{k}}t-2M\log^{2}t-\frac{4n}{a^{2}_{k-1}}-2t_{k-1}.$ (39)
In particular, for any $t\geq\frac{4}{\mu_{m_{k}w_{k}}}(2M\log^{2}\frac{4M}{\mu_{m_{k}w_{k}}}+\frac{2n}{a^{2}_{k-1}}+t_{k-1})$, we have
$Y^{t-1}_{m_{k}w_{k}}-Y^{t-1}_{m_{k}w}\geq\frac{\mu_{m_{k}w_{k}}}{2}t.$ (40)